Two object of masses 0.20 kg and 0.10 kg approach each other at 5.0 m/s and 2.5 m/s, respectively. Assuming that the collision is perfectly inelastic, what is the speed of the objects after the collision?
1
Expert's answer
2020-11-06T10:14:46-0500

According to the momentum conservation law, the equation of the momentums is:


"m_1v_1 + m_2v_2 = (m_1+ m_2)v'"

where "m_1 = 0.2kg", "m_2 = 0.1kg" are masses of the objects, "v_1 = 5m\/s", "v_2 = 2.5m\/s" are their speeds befor the collistion and "v'" is their speed after the collision. Since the collision is perfectly inelastic, both objects will move together with the same speed after the collision. Expressing "v'", obtain:


"v' = \dfrac{m_1v_1 + m_2v_2 }{m_1+ m_2} = \dfrac{0.2\times 5 + 0.1\times 2.5}{0.2+0.1} \approx 4.17m/s"

Answer. 4.17 m/s.